# Limits of Absoluteness of Observed Events in Timelike Scenarios: A No-Go Theorem

### v1

@@ Abstract
Wigner’s Friend–type paradoxes challenge the assumption that events are absolute — that when we measure a system, we obtain a single result, which is not relative to anything or anyone else. These paradoxes highlight the tension between quantum theory and our intuitions about reality being observer-independent. Building on a recent result that developed these paradoxes into a no-go theorem, namely the Local Friendliness Theorem, we introduce the Causal Friendliness Paradox, a time-ordered analogue of it. In this framework, we replace the usual locality assumption with Axiological Time Symmetry (ATS), and show that, when combined with the assumptions of Absoluteness of Observed Events (AOE), No Retrocausality (NRC), and Screening via Pseudo Events (SPE), we obtain a causal inequality. We then show that quantum mechanics violates this inequality and is therefore incompatible with at least one of these assumptions. To probe which assumption might be incompatible, we then examine whether AOE in its entirety is essential for this no-go result. We propose a weaker, operational form of AOE that still leads to inequalities which quantum mechanics violates. This result shows that even under relaxed assumptions, quantum theory resists reconciliation with classical notions of absolute events, reinforcing the foundational significance of Wigner’s Friend-type paradoxes in timelike scenarios.
@@ I Introduction
In Section III, we extend these results to a new paradox: the Causal Friendliness Paradox. In Section III.1, we describe the (hypothetical) experimental scenario for the Causal Friendliness Paradox to occur, before in Section III.2 giving the assumptions required for such a scenario to give predictions contradicting standard quantum mechanics (a contradiction we formally prove in Subsection III.3).
In Section IV, we expand on our result from Section III, by seeing to what extent we can weaken the Absoluteness of Observed Events assumption common to both our Causal Friendliness Paradox and the Extended Wigner’s Friend Paradoxes. We do this by looking at a result by Leifer and Pusey [17], where they weakened the $\psi$-ontic reproducibility assumption in Price’s 2007 no-go result on time-symmetry and retrocausality [16], to instead just require $\lambda$-mediation. In Section IV.1, we review how Leifer and Pusey weakened the $\psi$-ontic reproducibility assumption into $\lambda$-mediation. In Section IV.2, we then analogise this weakening at an operational, rather than ontological, level, to enable us to weaken the Absoluteness of Observed Events assumption while still being able to generate an inequality that quantum mechanics violates. However, in Section IV.3, we then show that we cannot weaken this assumption any further (i.e., specifically remove the Absoluteness of the Pseudo Events - the Friends measuring their system - which may or may not be unitarily “rewound” by the Wigners) without losing the ability to generate such a contradiction, with either quantum correlations or even Popescu and Rohrlich’s Boxworld’s correlations [18].
@@ II.2 Extended Wigner Friend Scenarios
Recently, a flurry of works has emerged [23, 24, 25, 26, 27, 28, 29], building upon the original Wigner’s Friend paradox [1] and its subsequent generalisations [30], and exploring various foundational issues related to these scenarios. Frauchiger and Renner [25] first proposed an extension of the Wigner’s Friend paradox in which the quantum probability assignments are shown to be logically inconsistent with the assumption of observer-independent facts. However, it was only in Brukner’s later version [23] that such inconsistencies were used to derive differing, testable predictions. Although these works received significant attention, their conclusions rely on counterfactual reasoning—an aspect which arguably weakened their claims. Bong et al later put forward a stronger no-go theorem [4], which, rather than challenging the notion of observer-independent facts involving counterfactual scenarios, directly questions the absoluteness of already observed events.
@@
It is important to note that, even though the events $c$ and $d$ are treated as observed events in Bong et al ’s treatment, this is not necessarily the case: when $x\neq 0$ ($y\neq 0$), event $c$ ($d$) is unitarily erased by Alice (Bob), as if a value was never assigned to the variable. Thus, we argue that events $a$,$b$ and events $c$,$d$ cannot be treated in the same way. When $x\neq 0$ ($y\neq 0$) we therefore call $c$ ($d$) Pseudo Events, while $a$ and $b$ we refer to as Truly-Observed Events.
Figure 2: Time-ordered measurement scenario for Causal Friendliness no-go theorem.
III Causal Friendliness Paradox
III.1 The Causal Friendliness Scenario
@@
III.2 The Causal Friendliness No-Go Theorem
Building upon our description of Truly-Observed Events and Pseudo Events, and inspired by Refs. [17, 16], we first define:
@@ III.2 The Causal Friendliness No-Go Theorem
A prepare-measure experiment with preparation $\mathbb{P}$ and measurement $\mathbb{M}$ has operational time symmetry if there exists a time-reversed experiment, with preparation $\mathbb{P}^{{}^{\prime}}$ and measurement $\mathbb{M}^{{}^{\prime}}$, in which $\mathbb{P}^{{}^{\prime}}\equiv\mathbb{M}$ and $\mathbb{M}^{{}^{\prime}}\equiv\mathbb{P}$ such that,
$p_{\leftarrow}(f,e|v,u)=p_{\rightarrow}(e,f|u,v)\ \forall u,v,x,y.$ (6)
where $e$ and $f$ are outcomes corresponding to observed events when the measurements $u$ and $v$ respectively are performed.
Less formally, if an experiment has the property of Operational Time Symmetry, this means that if the roles of preparations and measurements in the experiment are swapped, reversing their temporal order, the observed probabilities should remain unchanged. The temporal ordering of events is represented in the causal influence diagram shown in Fig. 3 for better understanding.
@@
Figure 3: Causal influence diagram for the Causal Friendliness scenario. The rectangular entities are those that are in direct control of the observer, while the circles represents events that are not in direct control. The arrows represents the order of the events conveying the cause-effect relationships between the events. Red coloured shapes indicate the Pseudo Events while the blue one are Truly-Observed Events.
@@
The impossibility of satisfying Eq. (8) is the main obstacle to having an Operationally Time Symmetric description of a theory. Nevertheless, without loss of generality, we may restrict our attention to those preparations for which Eq. (8) does hold. We refer to the corresponding set of correlations as no-signalling in time, denoted by $\mathcal{NS}_{T}$. It was shown in [17] that the sector of quantum theory satisfying $\mathcal{NS}_{T}$ is Operationally Time Symmetric. In what follows, while establishing our main results, we restrict to quantum preparations belonging to $\mathcal{NS}_{T}$ so that Operational Time Symmetry is guaranteed.
@@
It is important to note that the definition of Operational Time Symmetry pertains exclusively to events that are both real and operationally accessible. In contrast, the Causal Friendliness scenario includes not only these real events but also additional elements referred to as Pseudo Events. Consequently, in such a scenario, only an analogue of Operational Time Symmetry can be meaningfully defined. Motivated by this distinction, we employ this analogue to establish the main results of the paper.
@@
In Causal Friendliness scenario, the pair Charlie–Alice acts before Debbie–Bob in forward time, making the former the “preparers” and the latter the “measurers”. In the time-reversed picture, however, Debbie–Bob acts before Charlie–Alice, thus reversing the roles of “preparer” and “measurer”. At first glance, it may seem that as Charlie acts before Alice at a particular site of the whole scenario, this pair $(a,c)$ constitute a preparation and measurement in itself, where Charlie effectively prepares the system and Alice measures subsequently. Similar conclusions can be made for the other site where Debbie and Bob acts. However, this is not true. As already discussed, the events corresponding to $c$ and $d$ are Pseudo Events, as they are either eventually erased, or they end up corresponding to the outcomes $a$ and $b$ respectively. Therefore, we can only take the pair of events $(c,a)$ as a preparation, and the pair $(d,b)$ as a measurement, rather than splitting this up onto $a$ and $c$, and $b$ and $d$. This means the counterpart of the definition of Operational Time Symmetry applicable to this scenario does not involve reversing the entire temporal sequence $c\rightarrow a\rightarrow d\rightarrow b$ to $b\rightarrow d\rightarrow a\rightarrow c$. Instead, it reverses the temporal order of the grouped preparation and measurement events, namely $(c,a)\rightarrow(d,b)$, into $(d,b)\rightarrow(c,a)$. Considering this, a generalisation of Operational Time Symmetry can be proposed when this definition is extended to include the Pseudo Events. This forms the assumption of Axiological Time Symmetry, the final assumption needed for us to generate a full Causal Friendliness No-Go Theorem:
Theorem 1 (Causal Friendliness No-Go Theorem).
Consider the scenario we give in Section III.1, with Charlie and Alice acting before Debbie and Bob. Let $a,b,c,d\in\{\pm 1\}$ denote the outcomes of Alice, Bob, Charlie, and Debbie, respectively, and let $x,y\in\{0,1\}$ denote the measurement settings of Alice and Bob. Assume:
  1. 1.
    Absoluteness of Observed Events (AOE): All observed events at a particular point on the space-time region characterised by measurement outcomes $a,b,c,d$ have definite values, irrespective of any event that occurred at any other point in the space-time. This means the joint probability $p(c,a,d,b|x,y)$ exists and the observed probability $p(a,b|x,y)$ can be recovered from it via marginalisation as,
    $p(a,b|x,y)=\sum_{c,d}p(c,a,d,b|x,y)\ \ \forall x,y,a,b.$ (9)
  2. 2.
    Axiological Time Symmetry (ATS): If we invert the order of the scenario (i.e., have Debbie and Bob acting before Charlie and Alice), this inversion preserves the structure of these joint probabilities, i.e., $p_{\leftarrow}(d,b,c,a|y,x)=p_{\rightarrow}(c,a,d,b|x,y)\ \forall x,y,a,b,c,d$.
  3. 3.
@@
    No Retrocausality (NRC): Any future measurement choice $x(y)$ cannot influence the past outcomes corresponding to the Pseudo Events $c(d)$, and any observed event $a$ in the past cannot be influenced by the future Pseudo Event $d$.
  4. 4.
@@
    Screening via Pseudo Events (SPE): Every Truly-Observed Event is screened off from all the non-input variables in its causal past by appropriate pseudo-events.
    In the current scenario, for Bob’s observed event, this assumption implies:
@@
Given these Assumptions, the observed correlations satisfy the inequality:
@@
The variables $c$ and $d$, representing the Pseudo Events, impose constraints on Truly-Observed Events strictly weaker than those imposed by hidden variables in conventional no-go theorems. Hidden variables are typically assumed to encode complete information about a physical system and (at least in principle) enable an event-by-event description of its behaviour. In contrast, the SPE assumption endows $c$ and $d$ with only limited power, allowing them to influence solely those Truly-Observed Events lying within their future light cones, and screening the future events off from any past events that are already influenced by, or that are already influencing, these Pseudo Events. For example, if the Pseudo Event $c$ exists and is absolute, then the Truly-Observed Event $a$ must always depend on $c$, regardless of the choice of $x$. In particular, for $x=0$ one has $a=c$, while for other values of $x$, even if $c$ were erased and the laboratory restored to its prior state, the effect of any past information on $a$ could still be argued to be screened through the pre-existing value of $c$. In this sense, so long as there is the assumption (implicit in, but weaker than, AOE), that the Pseudo Events $c$ and $d$ are Absolute, then SPE emerges as a very natural assumption. We will discuss this implicit assumption further in Section IV.3, where we attempt to subdivide AOE to see which parts of it are actually necessary to generate a contradiction with quantum mechanics. While the implications of removing AOE as a whole have been considered elsewhere [33], we believe we are the first to consider the effect of sub-assumptions of AOE.
The retrocausality we appeal to with one of the parts of our NRC assumption takes a similar form to this retrocausality, except the prohibited relation is a causal dependence of the (potentially-observable) past Pseudo Events on the future measurement choices, rather than the dependence of any hidden variables on the measurement choices. This is as an instance of NRC (namely Pseudo Events are independent future measurement choices) where the Pseudo Events act equivalently to hidden variables in the conventional description of retrocausality. The other part of NRC comes from the idea that a Truly-Observed Past event $a$ cannot also depend on any of the non input variables $d$ that are potentially erased depending on a free choice.
From this point onward, we will collectively refer to the conjunction of the assumptions ATS, NRC, and SPE as Causal Agency. We now prove that the inequality given in Eq. (11) follows from the assumptions AOE and Causal Agency stated below.
III.3 Proof of the Causal Friendliness No-Go Theorem
Causal factorisation admitting NRC: In the forward time order, Charlie measures his quantum state first (obtaining $c$), then Alice chooses $x$ and so sets a value for $a$, then Debbie measures her quantum state (obtaining $d$), and finally Bob chooses $y$ and so sets a value for $b$. Thus, the most general causal factorisation for this scenario that admits NRC can be written as,

### v2

@@ Abstract
Wigner’s Friend–type paradoxes challenge the assumption that events are absolute — that when we measure a system, we obtain a single result, which is not relative to anything or anyone else. These paradoxes highlight the tension between quantum theory and our intuitions about reality being observer-independent. Building on a recent result that developed these paradoxes into a no-go theorem, namely the Local Friendliness Theorem, we introduce the Causal Time-Symmetric Friendliness Paradox, a time-ordered analogue of it. In this framework, we replace the usual locality assumption with Axiological Time Symmetry (ATS), and show that, when combined with the assumptions of Absoluteness of Observed Events (AOE), No Retrocausality (NRC), and Screening via Pseudo Events (SPE), we obtain a Causal Time-Symmetric Friendliness inequality. We then show that quantum mechanics violates this inequality and is therefore incompatible with at least one of these assumptions. To probe which assumption might be incompatible, we then examine whether AOE in its entirety is essential for this no-go result. We propose a weaker, operational form of AOE that still leads to inequalities which quantum mechanics violates. This result shows that even under relaxed assumptions, quantum theory resists reconciliation with classical notions of absolute events, reinforcing the foundational significance of Wigner’s Friend-type paradoxes in timelike scenarios.
@@ I Introduction
In Section III, we extend these results to a new paradox: the Causal Time-Symmetric Friendliness Paradox. In Section III.1, we describe the (hypothetical) experimental scenario for the Causal Time-Symmetric Friendliness Paradox to occur, before in Section III.2 giving the assumptions required for such a scenario to give predictions contradicting standard quantum mechanics (a contradiction we formally prove in Subsection III.3).
In Section IV, we expand on our result from Section III, by seeing to what extent we can weaken the Absoluteness of Observed Events assumption common to both our Causal Time-Symmetric Friendliness Paradox and the Extended Wigner’s Friend Paradoxes. We do this by looking at a result by Leifer and Pusey [39], where they weakened the $\psi$-ontic reproducibility assumption in Price’s 2007 no-go result on time-symmetry and retrocausality [48], to instead just require $\lambda$-mediation. In Section IV.1, we review how Leifer and Pusey weakened the $\psi$-ontic reproducibility assumption into $\lambda$-mediation. In Section IV.2, we then analogise this weakening at an operational, rather than ontological, level, to enable us to weaken the Absoluteness of Observed Events assumption while still being able to generate an inequality that quantum mechanics violates. However, in Section IV.3, we then show that we cannot weaken this assumption any further (i.e., specifically remove the Absoluteness of the Pseudo Events - the Friends measuring their system - which may or may not be unitarily “rewound” by the Wigners) without losing the ability to generate such a contradiction, with either quantum correlations or even Popescu and Rohrlich’s Boxworld’s correlations [47].
@@ II.2 Extended Wigner Friend Scenarios
Recently, a flurry of works has emerged [13, 14, 31, 23, 35, 50, 4, 69, 19], building upon the original Wigner’s Friend paradox [65] and its subsequent generalisations [20], and exploring various foundational issues related to these scenarios. Frauchiger and Renner [23] (potentially inspired by Brukner [13]) first proposed an extension of the Wigner’s Friend paradox in which the quantum probability assignments are shown to be logically inconsistent with the assumption of observer-independent facts. However, it was only in Brukner’s later version [14] that such inconsistencies were used to derive differing, testable predictions. Although these works received significant attention, their conclusions rely on counterfactual reasoning—an aspect which arguably weakened their claims. Bong et al later put forward a stronger no-go theorem [11], which, rather than challenging the notion of observer-independent facts involving counterfactual scenarios, directly questions the absoluteness of already observed events.
@@
It is important to note that, even though the events $c$ and $d$ are treated as observed events in Bong et al ’s treatment, this is not necessarily the case: when $x\neq 0$ ($y\neq 0$), event $c$, $d$ is unitarily erased by Alice (Bob), as if a value was never assigned to the variable. Thus, we argue that events $a$,$b$ and events $c$, $d$ cannot be treated in the same way. We therefore call $c$, $d$ Pseudo Events, while $a$ and $b$ we refer to as Truly-Observed Events.
We emphasise that the terms Pseudo Events and Truly-Observed Events should not be interpreted as universally- or ontologically-defined concepts. Rather, they are scenario-dependent notions introduced solely to specify the operational structure of the Causal Time-Symmetric Friendliness setup. More precisely, an event is labelled a Pseudo Event if, within the prescribed experimental protocol, there is ever the possibility of it being unitarily erased or being made operationally inaccessible by a super observer, while an event is labelled Truly-Observed event if it is never erased, and always remains empirically accessible in the scenario under consideration. This distinction is therefore relative to the chosen protocol and the level of control available to the (super)observers, and should not be interpreted as an absolute or fundamental classification of events. While it could be argued that, in quantum mechanics, all events are unitarily erasable, so all events are Pseudo Events, this has not been empirically demonstrated. Therefore, in the absence of any way to empirically demonstrate all events are unitarily erasable, we treat Truly-Observed Events as the rule and Pseudo Events as the exception, rather than the reverse.
Figure 2: Time-ordered measurement scenario for Causal Time-Symmetric Friendliness no-go theorem.
III Causal Time-Symmetric Friendliness Paradox
III.1 The Causal Time-Symmetric Friendliness Scenario
@@
We define a behaviour as the collection of joint probability distributions associated with all events in a given experimental scenario, including both Truly-Observed Events and Pseudo Events. A behaviour is specified by the set of joint probabilities $\mathcal{P}:=\{p(c,a,d,b|x,y)\}_{a,b,c,d,x,y}$ for all choices of inputs $x,y$, from which all operationally accessible marginals and conditional probabilities can be obtained by marginalisation. Further, we will call a behaviour $\mathcal{S}$ that satisfies certain specific properties a sector if $\mathcal{S}\subseteq\mathcal{P}$.
III.2 The Causal Time-Symmetric Friendliness No-Go Theorem
The main metaphysical assumption whose validity we aim to examine here is the same one used to establish the Local Friendliness inequality in [11]: the assumption of Absoluteness of Observed Events. However, the no-go theorem we intend to establish is formulated within the prepare-measure scenario discussed in the previous section and illustrated in Fig. 2. In this setting, the aforementioned assumption can be phrased as follows:
  1. 1.
    Absoluteness of Observed Events: All observed events at a particular point on the space-time region characterised by the measurement outcomes have definite values, irrespective of any event that occurred at any other point in the space-time.
    In Causal Time-Symmetric Friendliness scenario this implies for measurement events $a,b,c,d$ the joint probability $p(c,a,d,b|x,y)$ exists and the observed probability $p(a,b|x,y)$ can be recovered from it via marginalisation as,
    $p(a,b|x,y)=\sum_{c,d}p(c,a,d,b|x,y)\ \ \forall x,y,a,b.$ (6)
As discussed, the scenario we consider here can be cast as a conventional prepare-measure scenario consisting of a preparation stage $\mathbb{P}$ and a measurement stage $\mathbb{M}$. Each stage involves a specified set of inputs, which are directly controlled by the experimentalist, together with a corresponding set of outputs for each input, which are not directly controlled by the experimentalist. Building upon our description of observed events, and inspired by Refs. [39, 48], we first define:
@@ III.2 The Causal Time-Symmetric Friendliness No-Go Theorem
A theory satisfies operational time symmetry if for every prepare-measure experiment with preparation stage $\mathbb{P}$ and measurement stage $\mathbb{M}$, there exists a time-reversed experiment, with preparation stage $\mathbb{P}^{{}^{\prime}}$ and measurement stage $\mathbb{M}^{{}^{\prime}}$ (with same number of inputs and outputs as that of $\mathbb{M}$ and $\mathbb{P}$ respectively), such that joint probabilities of observed events remain the same, i.e.,
$p_{\leftarrow}(f,e|v,u)=p_{\rightarrow}(e,f|u,v)\ \forall u,v,e,f.$ (7)
[…]
Less formally, if an experiment has the property of Operational Time Symmetry, this means that if the roles of preparations and measurements in the experiment are swapped, reversing their temporal order, the observed probabilities should remain unchanged. The temporal ordering of events is represented in the causal influence diagram shown in Fig. 3 for better understanding. Clearly, the notion of time-symmetry adopted here, following [48, 39], differs from the conventional notion of time-symmetry typically discussed in the physics literature to express the symmetry of dynamics.
@@
Figure 3: Causal influence diagram for the Causal Time-Symmetric Friendliness scenario. The rectangular entities are those that are in direct control of the observer, while the circles represents events that are not in direct control. The arrows represents the order of the events conveying the cause-effect relationships between the events. Red coloured shapes indicate the Pseudo Events while the blue one are Truly-Observed Events.
@@
The impossibility of satisfying Eq. (9) is the main obstacle in having an Operationally Time Symmetric description of a theory. Nevertheless, without loss of generality, we may restrict our attention to those preparations for which Eq. (9) does hold. We refer to the corresponding set of behaviour as no-signalling in time, denoted by $\mathcal{NS}_{T}$. It was shown in [39] that the sector of quantum theory satisfying $\mathcal{NS}_{T}$ is Operationally Time Symmetric. In what follows, while establishing our main results, we restrict to quantum preparations belonging to $\mathcal{NS}_{T}$ so that Operational Time Symmetry is guaranteed.
@@
It is important to note that the definition of Operational Time Symmetry pertains exclusively to events that are both real and operationally accessible. In contrast, the Causal Time-Symmetric Friendliness scenario includes not only these real events but also additional elements referred to as Pseudo Events. Consequently, in such a scenario, only an analogue of Operational Time Symmetry can be meaningfully defined. Motivated by this distinction, we employ this analogue to establish the main results of the paper.
In the Causal Time-Symmetric Friendliness scenario, the pair Charlie-Alice acts before Debbie Bob in forward time, so the former pair naturally plays the role of the preparers while the latter acts as the measurers. In the time reversed description, however, Debbie-Bob acts before Charlie-Alice, thereby interchanging the roles of preparation and measurement. At first sight, one might therefore be tempted to interpret the local ordering at each site as defining an independent prepare-measure scenario, for instance by viewing Charlie as preparing the system and Alice as subsequently measuring it at one site, with an analogous interpretation for Debbie and Bob at the other site. However, such a description of the prepare-and-measure structure is ill-defined. To see this, consider a particular experimental run in which $x=1$ (or $y=1$), so that the corresponding event $c$ (or $d$) is erased and effectively treated as if it never occurred. In such a run, there is thus a measurement event without a corresponding preparation event, which shows that interpreting $c$ alone as a preparation is inconsistent. As discussed earlier, the events $c$ and $d$ are Pseudo Events, since they are either erased or ultimately mapped to the outcomes $a$ and $b$, respectively. Consequently, a consistent operational description only allows the pairs $(c,a)$ to be regarded as a preparation and $(d,b)$ as a measurement, rather than splitting these roles between the individual events $c$ and $a$, or $d$ and $b$.
@@
This means the counterpart of the definition of Operational Time Symmetry namely the Axiological Time Symmetry assumption applicable to this scenario does not involve reversing the entire temporal sequence $c\rightarrow a\rightarrow d\rightarrow b$ to $b\rightarrow d\rightarrow a\rightarrow c$. Instead, it reverses the temporal order of the grouped preparation and measurement events, namely $(c,a)\rightarrow(d,b)$, into $(d,b)\rightarrow(c,a)$. Considering this, a generalisation of Operational Time Symmetry can be proposed when this definition is extended to include the Pseudo Events. This forms the assumption of Axiological Time Symmetry, the final assumption needed for us to generate a full Causal Time-Symmetric Friendliness No-Go Theorem. Before stating the theorem, it is essential to formally define the assumptions and their implications, which are discussed in the following.
Definition 2 (Axiological Time Symmetry).
If the Operational Time Symmetry holds for Truly-Observed Events in a given scenario then it must also hold for all the events including the Pseudo Events involved in the scenario. In the taxonomy the Causal Time-Symmetric Friendliness settings this means that if statistics generated by the events corresponding to Alice’s and Bob’s measurements satisfies Operational Time Symmetry then in order to satisfy Axiological Time Symmetry the following relation must hold,
$p_{\leftarrow}(d,b,c,a|y,x)=p_{\rightarrow}(c,a,d,b|x,y)\ \forall x,y,a,b,c,d.$ (10)
We emphasise that we neither claim that Operational Time Symmetry holding at the level of Truly-Observed events means it must also hold for all events in scenarios involving Pseudo Events, nor that any underlying description incorporating pseudo events must satisfy Axiological Time Symmetry independently of the theory under consideration. Rather, in this work, we deliberately restrict our attention to cases in which Axiological Time Symmetry holds, and investigate what can be inferred about quantum theory under this assumption. Our aim in this paper is to examine the resulting constraints on observable correlations, assuming that Axiological Time Symmetry and the other stated assumptions hold, and to compare these constraints with quantum predictions. In the present scenario, Axiological Time Symmetry should be understood as an additional structural assumption extending Operational Time Symmetry beyond empirically accessible Truly-Observed Events to include Pseudo Events. While this extension is not compelled by operational considerations alone, we argue that it is a coherent and well motivated assumption within the specific Causal Time-Symmetric Friendliness framework considered here, and we make this restriction explicit while establishing our main results.
One might object the fact that the assumption of Axiological Time Symmetry already presupposes Absoluteness of Observed Events hold, since it is formulated in terms of a joint probability distribution over all events. We argue that although AOE is a prerequisite for defining ATS, this is a very natural and unavoidable pre-assumption rather than an additional hidden commitment. This point can be understood by analogy with Bell’s notion of local causality [7, 68]. In deriving a Bell inequality, one typically invokes realism, locality, and freedom of choice. However, locality cannot be meaningfully formulated without first specifying realism, as the notion of locality presupposes the existence of definite underlying events or variables to which locality can apply. In this sense, realism is already implicitly assumed when locality is defined. In an analogous, though conceptually distinct manner, it is still meaningful for us to investigate the effect of the assumption of Axiological Time Symmetry, even though it implicitly assumes Absoluteness of Observed Events. Specifically, in order to define Axiological Time Symmetry in probabilistic terms, one must first assume Absoluteness of Observed Events.
Definition 3 (No Retrocausality).
@@
Any future measurement choice cannot influence the past outcomes corresponding to the Pseudo Events and any Truly-observed event in the past cannot be influenced by the future Pseudo Events.
In the present scenario this assumption implies that events corresponding to the outcome $c(d)$ is independent of the choice of measurement $x(y)$ and the Truly-Observed Event $a$ is should be independent Pseudo Event $d$.
Here, the terms future and past are used in their standard physical sense - namely, an event $X$ is said to occur in the future (past) of another event $Y$ if it lies within the future (past) light cone of the spacetime region associated with $Y$. Moreover, throughout the paper, statements of the form “$X$ cannot influence YY”, in particular within the No Retrocausality assumption, are to be understood in a strictly operational and statistical sense. By this we mean that appropriate conditional independence relations hold for the relevant probability distributions, for example
$\begin{split}&p(c|x,y)=p(c),\\ &p(d|y)=p(d),\\ &p(a|c,x,d)=p(a|c,x).\end{split}$ (11)
We do not assume a general causal Markov condition associated with directed acyclic graph-based causal models [45]. Instead, these conditional independences are postulated directly as scenario-specific restrictions motivated by the operational structure of the Causal Time-Symmetric Friendliness setup. The phrase “cannot influence” should therefore be read exclusively as a shorthand for these operational conditional independence constraints.
The retrocausality we appeal to with one of the parts of our No Retrocausality assumption takes a similar form to this retrocausality, except the prohibited relation is a causal dependence of the (potentially-observable) past Pseudo Events on the future measurement choices, rather than the dependence of any hidden variables on the measurement choices. This is as an instance of No Retrocausality (namely Pseudo Events are independent future measurement choices) where the Pseudo Events act equivalently to hidden variables in the conventional description of retrocausality. The other part of No Retrocausality comes from the idea that a Truly-Observed Past event $a$ cannot also depend on any of the non input variables $d$ that are potentially erased depending on a free choice.
Note that, in no part of the paper do we claim that $c,d$ should be regarded as something that exists independent of the observation that contains intrinsic properties of the system. Rather, we say if a Pseudo Event $c$ or $d$ occurs then it creates accessible information about the system (even if erased) which can in principle influence any future event corresponding to any future measurement outcome. This inspires us to define the final metaphysical assumption as follows:
Definition 4 (Screening via Pseudo Events).
@@
Every Truly-Observed Event is screened off from all the non-input variables that lies in its causal past by Pseudo Events that its past light cone contains. In the current scenario, for Bob’s observed event, this assumption implies:
@@
The variables $c$ and $d$, representing the Pseudo Events, impose constraints on Truly-Observed Events strictly weaker than those imposed by hidden variables in conventional no-go theorems. Hidden variables are typically assumed to encode complete information about a physical system and (at least in principle) enable an event-by-event description of its behaviour. In contrast, the Screening via Pseudo Events assumption endows $c$ and $d$ with only limited power, allowing them to influence solely those Truly-Observed Events lying within their future light cones, and screening the future events off from any past events that are already influenced by, or that are already influencing, these Pseudo Events. For example, if the Pseudo Event $c$ exists and is absolute, then the Truly-Observed Event $a$ must always depend on $c$, regardless of the choice of $x$. In particular, for $x=0$ one has $a=c$, while for other values of $x$, even if $c$ were erased and the laboratory restored to its prior state, the effect of any past information on $a$ could still be argued to be screened through the pre-existing value of $c$. In this sense, so long as there is the assumption (implicit in, but weaker than, Absoluteness of Observed Events), that the Pseudo Events $c$ and $d$ are Absolute, then Screening via Pseudo Events emerges as a very natural assumption. We will discuss this implicit assumption further in Section IV.3, where we attempt to subdivide Absoluteness of Observed Events to see which parts of it are actually necessary to generate a contradiction with quantum mechanics. While the implications of relaxing Absoluteness of Observed Events as a whole have been considered in literature [3, 60, 67], we believe we are the first to consider the effect of sub-assumptions of Absoluteness of Observed Events.
Theorem 1 (Causal Time-Symmetric Friendliness No-Go Theorem).
Consider the scenario we give in Section III.1, with Charlie and Alice acting before Debbie and Bob. Let $a,b,c,d\in\{\pm 1\}$ denote the outcomes of Alice, Bob, Charlie, and Debbie, respectively, and let $x,y\in\{0,1\}$ denote the measurement settings of Alice and Bob. If the operational statistics admits Absoluteness of Observed Events, Axiological Time Symmetry, No Retrocausality, and Screening via Pseudo Events , the observed correlations satisfy the inequality:
@@
From this point onward, we will collectively refer to the conjunction of the metaphysical assumptions Axiological Time Symmetry, No Retrocausality, and Screening via Pseudo Events as Causal Agency. We now prove that the inequality given in Eq. (13) follows from the assumptions Absoluteness of Observed Events and Causal Agency stated below.
III.3 Proof of the Causal Time-Symmetric Friendliness No-Go Theorem
Causal factorisation admitting No Retrocausality: In the forward time order, Charlie measures his quantum state first (obtaining $c$), then Alice chooses $x$ and so sets a value for $a$, then Debbie measures her quantum state (obtaining $d$), and finally Bob chooses $y$ and so sets a value for $b$. Thus, the most general causal factorisation for this scenario that admits No Retrocausality can be written as,